# Completely entangled subspaces of entanglement depth $k$

### v2

@@ V.1 Bound on cardinalities of $\mathrm{UPB}_{\mathrm{k-1}}$
The trivial bounds from Eq. (25) can in many cases be strengthened if the internal product structure of partitions is appropriately taken into account. The improvement can be achieved by a generalization of the technique recently introduced in Ref. [57] to lower-bound permissible cardinalities of UPBn-1, i.e., GUPBs. There, the pigeonhole principle was the key resource, here, its generalization comes naturally in handy. We have the following.
@@
We will show that if the number of elements in $\mathcal{K}$ is strictly smaller than the RHS of (26) then $\mathcal{K}$ is extendible with vectors of the entanglement depth $k-1$ or smaller, i.e., it is not a $\mathrm{UPB}_{\mathrm{k-1}}$.
@@
Let us focus on one of the vectors from $\mathcal{K}$, e.g., $|v_{1}\rangle$. Assume that the said $s$ vectors which are orthogonal to $|v_{1}\rangle$ are $\mathcal{K}_{1}=\{|v_{2}\rangle,|v_{3}\rangle,\ldots,|v_{s+1}\rangle\}$ with orthogonality holding on sites from $\textbf{A}_{w}:=A_{1},A_{2},\cdots,A_{w}$. Since there exist a vector orthogonal to $\mathcal{K}_{1}$ on one of the sites from $\textbf{A}_{w}$, one immediately infers that the elements of $\mathcal{K}_{1}$ do not span the whole Hilbert space corresponding to $\textbf{A}_{w}$, that is
@@
Now, if the remaining vectors $\mathcal{K}_{2}:=\mathcal{K}\setminus\mathcal{K}_{1}=\{|v_{1}\rangle,|v_{s+2}\rangle,\ldots,|v_{m}\rangle\}$ did not span on $\mathbf{A}\setminus\textbf{A}_{w}$ the corresponding Hilbert space, i.e.,
@@
then one would find a vector orthogonal to $\mathcal{K}$ in the form
@@
with an arbitrary $(n-w)$-partite vector $|\xi\rangle$ orthogonal to $\mathcal{K}_{2}^{\mathbf{A}\setminus\textbf{A}_{w}}$. Clearly, the vector $|\xi\rangle$ would have in this case the entanglement depth at most $(n-w)$ and in consequence $\mathcal{K}$ would not be a $\mathrm{UPB}_{\mathrm{n-w}}$.
@@
Let us define the following function corresponding to the LHS of Eq. (34)
@@
We now observe that we can replace the maximum with the function inside of it, as for any $m$ for which this function is less than one, we can find another argument $\tilde{m}>m$ attaining one without changing the value of the whole function under scrutiny; precisely, we can take $\tilde{m}=n\lfloor\frac{m-1}{n}\rfloor+w+1$ for which $f_{w}(m)=f_{w}(\tilde{m})$. In turn, we can consider
@@
Plugging back $r=d^{n-w}-1$ and substituting $w=n-k+1$ [(33)], we finally obtain the largest excluded cardinality
@@
We have proved the statement for equal local dimension but a similar reasoning can be applied to the general case. We omit the derivation here.
We now show when the obtained bound is non-trivial, i.e., it gives a lower bound strictly larger than the one provided by Eq. (25). We have the following chain of inequalities:
@@
Note that for $k=n$ (GUPBs) the bound reproduces, as it should, the one from [57], in which case it is always non-trivial. This bound has been recently improved in [59]. At this point, it is not clear whether our current bound on the permissible cardinalities of $\mathrm{UPB}_{\mathrm{k-1}}$’s can be improved using the graph theory techniques of [59].
@@ V.2 Construction of $k$-CESs from UPBs
While the bound of Proposition 5 puts limitations on sizes of $\mathrm{UPB_{k-1}}$’s, it does not say anything about the possibility of the actual constructions of bases with permissible cardinalities. This is particularly important in view of the unknown status of this problem for GUPBs ($k=n$). It turns out that $\mathrm{UPB_{k-1}}$’s do exist for $k<n$ and in fact there are already examples in the literature for systems of four qubits [60].
We will now reconstruct one of the four-qubit basis from Ref. [60] and show existence of a $\mathrm{UPB_{3}}$ in the case of five qubits. In both situations, we are interested in the case of $k=n-1$. With this aim we use the results of Ref. [61], where a recursive procedure was given for a construction of UPBs of size $2^{n-1}$ for any $n$.
@@
which recovers one of the UPBs given in Ref. [60], where it was shown that it is unextendible for any of the two vs two parties bipartitions, i.e., $\overline{\mathcal{K}^{(4)}}$ is a $\mathrm{UPB_{2}}$. As noted in Section V.1, its cardinality is the minimal possible.
@@
The set has a visibly cyclic structure and when proving (non)existence of a vector with the entanglement depth three orthogonal to all its elements one needs to consider only two bipartitions as the rest are equivalent to one of them. These bipartitions are $A_{1}A_{2}|A_{3}A_{4}A_{5}$ and $A_{1}A_{3}|A_{2}A_{4}A_{5}$. With an exhaustive search based on Fact 1 we have verified that $\mathcal{K}^{(5)}$ is already a $\mathrm{UPB}_{3}$ and no modifications of it are necessary.
Since the UPBs of Ref. [61] are constructed recursively, it is plausible that they can be used, directly or after some modifications similar to those shown above in the four qubit case, to obtain $\mathrm{UPB_{n-2}}$’s for any $n$.
Let us conclude with an observation that in general the case $k-1|n$ is somewhat special since then we deal with systems with $t$ subsystems of equal local dimensions $d^{k-1}$ (recall our assumptions of dimension $d$ of each subsystem $A_{i}$) in the ”coarse-grained” Hilbert space [Eq. (14)]. Constructions of UPBs for homogeneous systems (i.e., with equal local dimensions) are far better explored that those for heterogeneous systems (i.e., different local dimensions). A possible route to find $\mathrm{UPB}_{\mathrm{k-1}}$’s could then be to build UPBs first for (14) and then look for their proper ”fine-grained” versions in the original space.
@@ VI Conclusions and outlook
We have proposed to consider the notion of the entanglement depth in the context of entangled subspaces by introducing the notion of completely entangled subspaces of entanglement depth $k$ ($k$-CESs), that is subspaces composed solely of pure states with the entanglement depth at least $k$. We have presented a universal construction of such subspaces working for any multipartite systems. We have also considered the relationship between unextendible product bases (UPBs) and $k$-CESs. In particular, we have provided a non-trivial bound on the cardinalities of UPBs whose orthocomplements are $k$-CESs. Further, we have discussed the problem of constructing such UPBs and provided their examples in systems of several qubits.
From the general point of view, it is natural to expect that the newly introduced notion may turn out relevant for the analyses of many-body systems, where the entanglement depth is a figure of merit.  The already established connection between entangled subspaces and quantum error correction may suggest that $k$-CESs could be of some use also in this area and their analysis could help understand this relationship deeper. Moreover, it would be interesting to see whether the $k$-CESs we have provided could be useful in further studies on counterexamples to the additivity of the minimum output Renéyi entropy of quantum channels.
There are also several concrete open problems naturally emerging from our study. For example, it would be desirable to see if the obtained bound on the cardinalities of UPBs leading to $k$-CESs can be further improved, or, to research the possibility of a general practical construction of such UPBs, as this would automatically entail a construction of PPT states. Finally, there is constant demand for practical methods of subspace entanglement certification and quantification tailored to specific types of entangled subspaces and this line or research is also worth exploration.
@@ VII Acknowledgments
K.V. and R. A. acknowledge the support of the National Science Center (Poland) through the SONATA BIS project (grant no. 2019/34/E/ST2/00369). Enlightening discussions with Marcin Wieśniak are acknowledged.
@@ Appendix A Minimal value of (12)
Let $n$ and $k\leq n$ be positive integers. Let further
@@
(see, e.g., [63]) Let $g(\cdot)$ be a convex function and $\boldsymbol{x}\prec\boldsymbol{y}$. It then holds
@@ Appendix B Properties of $D_{k-CES}^{max}$
Case (a). Assume now $d$ to be a continuous parameter such that $d\in[2,\infty)$. It holds
@@
(notice that $t_{n,k}\geq t_{n,{k+1}}$ and we may need to pad zeros to $\tilde{\mathbb{y}}$ so that both vector match in length). Clearly, $\mathbb{y}\prec\tilde{\mathbb{y}}$ and by Lemma 2 it holds $D(d,k+1,n)<D(d,k,n)$. $\square$
@@
Since $2\leq k\leq n$ one easily verifies that $D(d,k,n+1)>D(d,k,n)$. $\square$
@@ Appendix C Maximal $3$-CES: four qubit case
In this appendix, we give a nine-dimensional, i.e., maximal, $3$-CES in the case of four qubits ($n=4$, $d=2$, $k=3$) obtained with the construction of Section IV.
With the choice of the nodes $x_{i}=i+1$, $i=0,1,\dots,6$, the Vandermonde vectors (21) are now of the form (we omit the transpose and shift the index)
@@
$(1,i^{8})_{A_{1}}\otimes(1,i^{4})_{A_{2}}\otimes(1,i^{2})_{A_{3}}\otimes(1,i)_{A_{4}}$ (67)
$i=1,2,\dots,7$. The matrix built from these vectors is totally positive. This implies that matrices constructed from bipartite vectors on any pair $A_{m}A_{n}$ [cf. Eq. 24] are always full rank, meaning, in particular, that any four ($d^{2}$ with $d=2$) such vectors span the whole four dimensional Hilbert space. Corollary 1 tells us then that there does not exist a vector of the form $|\phi\rangle_{A_{m}A_{n}}\otimes|\varphi\rangle_{\mathbb{A}\setminus A_{m}A_{n}}$, i.e., a $2$-producible vector, which is orthogonal to all the vectors (C). Consequently, the orthocomplement of the span of the vectors (C) is a $3$-CES.
@@ Appendix D Generalized pigeonhole principle
Let us start with recalling the definition of the function in question:
@@
Subcase (ii,a): $m-n\displaystyle\left\lfloor\frac{m}{n}\right\rfloor-w\geq 1$. We just note that this condition can never be satisfied simultaneously with case (i) and is only valid for $n\nmid m$. We have
@@
in view of Eq. (84). In consequence, in both $f_{w}(m)$ and $f_{w}(m+1)$ we can take the function under the maximum as its value. We then have after trivial simplifications

### v3

@@ V.1 Bound on cardinalities of $\mathrm{UPB}_{\mathrm{k-1}}$
The trivial bounds from Eq. (25) can in many cases be strengthened if the internal product structure of partitions is appropriately taken into account. The improvement can be achieved by a generalization of the technique recently introduced in Ref. [57] to lower-bound permissible cardinalities of UPBn-1, i.e., GUPBs. There, the pigeonhole principle was the key resource, here, its generalization naturally comes in handy. We have the following.
@@
We will show that if the number of elements in $\mathcal{K}$ is strictly smaller than the right-hand side of (26) then $\mathcal{K}$ is extendible with vectors of the entanglement depth $k-1$ or smaller, i.e., it is not a $\mathrm{UPB}_{\mathrm{k-1}}$.
@@
Let us focus on one of the vectors from $\mathcal{K}$, e.g., $|v_{1}\rangle$. Assume that the said $s$ vectors which are orthogonal to $|v_{1}\rangle$ are $\mathcal{K}_{1}=\{|v_{2}\rangle,|v_{3}\rangle,\ldots,|v_{s+1}\rangle\}$ with orthogonality holding on sites from $\textbf{A}_{w}:=A_{1},A_{2},\cdots,A_{w}$. Since there exist a vector orthogonal to $\mathcal{K}_{1}$ on one of the sites from $\textbf{A}_{w}$, we immediately infer that the elements of $\mathcal{K}_{1}$ do not span the whole Hilbert space corresponding to $\textbf{A}_{w}$, that is
@@
Now, if the remaining vectors $\mathcal{K}_{2}:=\mathcal{K}\setminus\mathcal{K}_{1}=\{|v_{1}\rangle,|v_{s+2}\rangle,\ldots,|v_{m}\rangle\}$ do not span on $\mathbf{A}\setminus\textbf{A}_{w}$ the corresponding Hilbert space, i.e.,
@@
then we will find a vector orthogonal to $\mathcal{K}$ in the form
@@
with an arbitrary $(n-w)$-partite vector $|\xi\rangle$ orthogonal to $\mathcal{K}_{2}^{\mathbf{A}\setminus\textbf{A}_{w}}$. Clearly, the vector $|\xi\rangle$ would have, in this case, an entanglement depth of at most $(n-w)$, and as a consequence, $\mathcal{K}$ would not be a $\mathrm{UPB}_{\mathrm{n-w}}$.
@@
Let us define the following function corresponding to the left-hand side of Eq. (34)
@@
We now observe that we can replace the maximum with the function inside of it, as for any $m$ for which this function is less than one, we can find another argument $\tilde{m}>m$ that reaches $1$ without changing the value of the whole function under scrutiny; precisely, we can take $\tilde{m}=n\lfloor\frac{m-1}{n}\rfloor+w+1$ for which $f_{w}(m)=f_{w}(\tilde{m})$. In turn, we can consider
@@
Plugging $r=d^{n-w}-1$ back in and substituting $w=n-k+1$ [(33)], we finally obtain the largest excluded cardinality
@@
We have proved the statement for equal local dimensions but a similar reasoning can be applied to the general case. We omit the derivation here.
We now show when the obtained bound is non-trivial, i.e., when it gives a lower bound strictly larger than the one provided by Eq. (25). We have the following chain of inequalities:
@@
Note that for $k=n$ (GUPBs) the bound reproduces, as it should, the one from [57], in which case it is always non-trivial. This bound was recently improved in [59]. At this point, it is not clear whether our current bound on the permissible cardinalities of $\mathrm{UPB}_{\mathrm{k-1}}$’s can be improved using the graph theory techniques of [59].
@@ V.2 Construction of $k$-CESs from UPBs
While the bound of Proposition 5 puts limitations on the sizes of $\mathrm{UPB_{k-1}}$’s, it does not say anything about the possibility of the actual constructions of bases with permissible cardinalities. This is particularly important in view of the unknown status of this problem for GUPBs ($k=n$). It turns out that $\mathrm{UPB_{k-1}}$’s do exist for $k<n$ and in fact there are already examples in the literature for systems of four qubits [60].
We will now reconstruct one of the four-qubit basis from Ref. [60] and show the existence of a $\mathrm{UPB_{3}}$ in the case of five qubits. In both situations, we are interested in the case of $k=n-1$. For this aim we use the results of Ref. [61], where a recursive procedure was given for a construction of UPBs of size $2^{n-1}$ for any $n$.
@@
which recovers one of the UPBs given in Ref. [60], which showed that it is unextendible for any of the two vs two party bipartitions, i.e., that $\overline{\mathcal{K}^{(4)}}$ is a $\mathrm{UPB_{2}}$. As noted in Section V.1, its cardinality is the minimal possible.
@@
The set has a visibly cyclic structure and when proving the (non)existence of a vector with an entanglement depth of $3$ orthogonal to all its elements one needs to consider only two bipartitions because the rest are equivalent to one of them. These bipartitions are $A_{1}A_{2}|A_{3}A_{4}A_{5}$ and $A_{1}A_{3}|A_{2}A_{4}A_{5}$. With an exhaustive search based on Fact 1 we have verified that $\mathcal{K}^{(5)}$ is already a $\mathrm{UPB}_{3}$ and no modifications of it are necessary.
Since the UPBs of Ref. [61] were constructed recursively, it is plausible that they could be used, directly or after some modifications similar to those shown above in the four qubit case, to obtain $\mathrm{UPB_{n-2}}$’s for any $n$.
Let us conclude with an observation that in general the case $k-1|n$ is somewhat special because then we deal with systems with $t$ subsystems of equal local dimensions $d^{k-1}$ (recall our assumptions of dimension $d$ of each subsystem $A_{i}$) in the ”coarse-grained” Hilbert space [Eq. (14)]. Constructions of UPBs for homogeneous systems (i.e., with equal local dimensions) are far better explored that those for heterogeneous systems (i.e., different local dimensions). A possible route to find $\mathrm{UPB}_{\mathrm{k-1}}$’s could then be to build UPBs first for (14) and then look for their proper ”fine-grained” versions in the original space.
@@ VI Conclusions and outlook
We proposed to consider the notion of the entanglement depth in the context of entangled subspaces by introducing the notion of completely entangled subspaces of entanglement depth $k$ ($k$-CESs), that is subspaces composed solely of pure states with an entanglement depth of at least $k$. We presented a universal construction of such subspaces that works for any multipartite systems. We also considered the relationship between unextendible product bases (UPBs) and $k$-CESs. In particular, we provided a non-trivial bound on the cardinalities of UPBs whose orthocomplements are $k$-CESs. Further, we discussed the problem of constructing such UPBs and provided some examples in systems of several qubits.
From the general point of view, it is natural to expect that the introduced notion may turn out to be relevant for analyses of many-body systems, where the entanglement depth is a figure of merit. The already established connection between entangled subspaces and quantum error correction may suggest that $k$-CESs could be of some use also in this area and their analysis could help us to understand this relationship better. Moreover, it would be interesting to see whether the $k$-CESs we have provided could be useful in further studies on counterexamples to the additivity of the minimum output Rényi entropy of quantum channels.
There are also several concrete open problems that naturally emerge from our study. For example, it would be desirable to see whether the obtained bound on the cardinalities of UPBs leading to $k$-CESs can be further improved, in particular, using graph-theoretic methods, and to research the possibility of a general practical construction of such UPBs, as this would automatically entail a construction of PPT states. Finally, there is constant demand for practical methods of subspace entanglement certification and quantification tailored to specific types of entangled subspaces and this line or research is also worth exploration.
@@ VII Acknowledgments
K.V. and R. A. acknowledge the support of the National Science Center (Poland) through the SONATA BIS project (grant no. 2019/34/E/ST2/00369). Enlightening discussions with M. Wieśniak are acknowledged.
@@ Appendix A Minimal value of (12)
Let $n$ and $k\leq n$ be positive integers. Further, let
@@
(see, e.g., [63]) Let $g(\cdot)$ be a convex function and $\boldsymbol{x}\prec\boldsymbol{y}$. It then holds that
@@ Appendix B Properties of $D_{k-CES}^{max}$
Case (a). Assume now $d$ is a continuous parameter such that $d\in[2,\infty)$. It holds
@@
(notice that $t_{n,k}\geq t_{n,{k+1}}$ and we may need to pad zeros in $\tilde{\mathbb{y}}$ so that the vectors match in length). Clearly, $\mathbb{y}\prec\tilde{\mathbb{y}}$ and by Lemma 2 it holds that $D(d,k+1,n)<D(d,k,n)$. $\square$
@@
Since $2\leq k\leq n$ we can easily verify that $D(d,k,n+1)>D(d,k,n)$. $\square$
@@ Appendix C Maximal $3$-CES: four qubit case
In this appendix, we give a nine-dimensional, i.e., maximal, $3$-CES in the case of four qubits ($n=4$, $d=2$, $k=3$) obtained with the construction in Sec. IV.
With the choice of nodes $x_{i}=i+1$, $i=0,1,\dots,6$, the Vandermonde vectors (21) are now of the form (we omit the transpose and shift the index)
@@
$(1,i^{8})_{A_{1}}\otimes(1,i^{4})_{A_{2}}\otimes(1,i^{2})_{A_{3}}\otimes(1,i)_{A_{4}},$ (67)
with $i=1,2,\dots,7$. The matrix built from these vectors is totally positive. This implies that matrices constructed from bipartite vectors on any pair $A_{m}A_{n}$ [cf. Eq. 24] are always full rank, meaning, in particular, that any four ($d^{2}$ with $d=2$) such vectors span the whole four dimensional Hilbert space. Corollary 1 tells us then that there does not exist a vector of the form $|\phi\rangle_{A_{m}A_{n}}\otimes|\varphi\rangle_{\mathbb{A}\setminus A_{m}A_{n}}$, i.e., a $2$-producible vector, which is orthogonal to all the vectors (C). Consequently, the orthocomplement of the span of the vectors (C) is a $3$-CES.
@@ Appendix D Generalized pigeonhole principle
Let us start by recalling the definition of the function in question:
@@
Subcase (ii,a): $m-n\displaystyle\left\lfloor\frac{m}{n}\right\rfloor-w\geq 1$. We just note that this condition can never be satisfied simultaneously with case (i) and is valid only for $n\nmid m$. We have
@@
in view of Eq. (84). As a consequence, in both $f_{w}(m)$ and $f_{w}(m+1)$ we can take the function under the maximum as its value. We then have after trivial simplifications